OpenStudy (anonymous):

plz help with Your friend hands you a graph of the performance statistics of the newest car to roll off the assembly line. He says, “I know this graph is f(x) = –4(x – 5)3 + 7 but I can’t remember how it is related to the graph of x3.” Explain to your friend how the graph f(x) is a translation of the graph x3.

OpenStudy (anonymous):

y=(x)3 Let's try to turn this into the performance function. Putting a negative in front flips it over the x-axis. y=−(x)3 To make shifts to the right or up we subtract that value from the corresponding variable. So we want to shift this function to the right 5 units, so we'll subtract 5 from x. y=−(x−5)3 We also want to shift this function up by 7, so we'll subtract 7 from y. y−7=−(x−5)3 Then add 7 to each side, y=−(x−5)3+7 Then we'll squeeze the function (make it grow four times faster) by multiplying it by 4. y=−4(x−5)3+7 via http://openstudy.com/study#/updates/53a1d53ae4b024ecf3c51436